OpenStudy (anonymous):

Tamara gets paid an annual salary of $48,000 and earns 4% commission on all sales she makes. If Tamara wants to make $6,000 this month, how much in sales does she need to have?

jimthompson5910 (jim_thompson5910):

if she makes x dollars in sales, then how much commission does she get?

OpenStudy (anonymous):

I don't know, that's all the question says

jimthompson5910 (jim_thompson5910):

what is 4% in decimal form

OpenStudy (anonymous):

.04

jimthompson5910 (jim_thompson5910):

so if she makes x dollars in sales, then she makes 0.04x dollars in commission

jimthompson5910 (jim_thompson5910):

her annual salary is $48,000 that makes her monthly salary: 48000/12 = 4000 dollars

jimthompson5910 (jim_thompson5910):

add the commission to her salary to get 0.04x+4000

jimthompson5910 (jim_thompson5910):

She's making 0.04x+4000 dollars per month (x is the amount in sales)

jimthompson5910 (jim_thompson5910):

She wants to make $6,000 this month, so 0.04x + 4000 = 6000 solve for x to get your final answer

OpenStudy (anonymous):

Still not getting it. Just gonna go with $50,000, thanks anyway.

jimthompson5910 (jim_thompson5910):

do you see how I got 0.04x + 4000

OpenStudy (anonymous):

Yeah, the 0.04 is her commission, x is the amount in sales and 4000 is her monthly salary

jimthompson5910 (jim_thompson5910):

so we have commission + salary = 0.04x + 4000

jimthompson5910 (jim_thompson5910):

we want that equal to 6000 dollars, so that's why 0.04x + 4000 = 6000

jimthompson5910 (jim_thompson5910):

Now solve for x 0.04x + 4000 = 6000 0.04x = 6000 - 4000 0.04x = 2000 x = 2000/0.04 x = 50,000

jimthompson5910 (jim_thompson5910):

So she has to make $50,000 in sales (ie sell a bunch of products that add up to $50,000 in value) to earn $6,000 this month

OpenStudy (anonymous):

Okay, I get it now, had to write it out haha I have a similar problem I did the same process with but got an answer that's not an option. Bob gets paid an annual salary of $30,000 and earns 5% commission on all sales he makes. If Bob wants to make $6,000 this month, how much in sales does he need to have? a.$50,000 b.$70,000 c.$90,000 d.$120,000

OpenStudy (anonymous):

I'm getting 80,000

jimthompson5910 (jim_thompson5910):

what equation did you set up?

OpenStudy (anonymous):

I just found where I messed up writing it out, thanks again!

jimthompson5910 (jim_thompson5910):

what did you get this time?

OpenStudy (anonymous):

70000

jimthompson5910 (jim_thompson5910):

very good